### v1

# Elimination Without Eliminating: Computing Complements of Real Hypersurfaces Using Pseudo-Witness Sets

Abstract
Many hypersurfaces in algebraic geometry, such as discriminants, arise as the projection of another variety. The real complement of such a hypersurface partitions its ambient space into open regions. In this paper, we propose a new method for computing these regions. Existing methods for computing regions require the explicit equation of the hypersurface as input. However, computing this equation by elimination can be computationally demanding or even infeasible. Our approach instead derives from univariate interpolation by computing the intersection of the hypersurface with a line. Such an intersection can be done using so-called pseudo-witness sets without computing a defining equation for the hypersurface – we perform elimination without actually eliminating. We implement our approach in a forthcoming Julia package and demonstrate, on several examples, that the resulting algorithm accurately recovers all regions of the real complement of a hypersurface.
Keywords. hypersurface complement, routing function, pseudo-witness set, real numerical algebraic geometry, homotopy continuation, numerical algebraic geometry
2020 AMS Subject Classification. 65H10, 14Q20, 14P25, 68W30.
1 Introduction
Many hypersurfaces in algebraic geometry arise as the Zariski closure of the projection of another variety. The real complement of such a hypersurface partitions its real ambient space into open regions. This paper presents a new method for computing these regions without requiring a defining equation for the hypersurface. By computing these regions, we mean representing them with a finite set of sample points from which one can determine the number of regions, do membership tests, and extract topological information such as Euler characteristics, for example.
Several algorithms have been developed to compute the regions of a hypersurface $\mathcal{H}$}, assuming an explicit defining polynomial $h$ for $\mathcal{H}$}; i.e., ${\mathcal{H}}=V(h)\subset{\mathbb{C}}^{n}$. See for example [3, Section 15] and [12, 46]. One such approach [11, 17, 36, 35] uses so-called routing functions of the form
$r(x)=\frac{|h(x)|}{(1+\lVert x-c\rVert^{2})^{e}}=\frac{|h(x)|}{(1+(x_{1}-c_{1})^{2}+\cdots+(x_{n}-c_{n})^{2})^{e}},$ (1.1)
where $c\in{\mathbb{R}}^{n}$ is generic and $e$ is a positive integer with $2e>\deg h$. This ensures that $r$ vanishes on ${\mathcal{H}}\cap{\mathbb{R}}^{n}$ and at infinity, so that $r$ has at least one critical point in every region of ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$. Computing these critical points and connecting them via gradient flow of $r$ yields a graph whose connected components represent the regions of ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$. This approach has found applications in, for instance, convex geometry [9] or ecology [16, 13], and has been implemented in the Julia package HypersurfaceRegions.jl [11].
One challenge in using routing functions in applications is that many hypersurfaces of interest arise through projection. In this case, obtaining an explicit defining polynomial $h$ of the hypersurface requires a computationally expensive symbolic elimination, which even for moderate problems is often out of reach with current technology, and thus prohibits the direct use of routing functions.
One of the main observations of this work is that the following tasks can be carried out without explicitly computing the polynomial $h$:
  1. (1)
    compute the degree $\deg h$;
  2. (2)
    evaluate $h(x)$ at points $x\in\mathbb{C}^{n}$ up to scaling;
  3. (3)
    evaluate the gradient $\nabla h(x)$ at points $x\in\mathbb{C}^{n}$; and
  4. (4)
    evaluate the Hessian ${\rm hess}(h)(x)$ at points $x\in\mathbb{C}^{n}$.
These operations are enough to define a routing function $r$, run homotopy continuation [4, 5, 50] on $\nabla r$ to determine critical points, and compute gradient flow of $r$ (in fact, in Section 3 we will work with $\log r$ instead). The key idea is to replace the explicit defining polynomial $h$ with numerical data obtained from a pseudo-witness set [30] that represents $\mathcal{H}$} through its intersection with a generic line.
Pseudo-witness sets have previously been used for membership tests [23], finding generators of certain elimination ideals [24], and, more generally, for numerical implicitization [14]. Recently, in [22], it was proposed that pseudo-witness sets can be used for evaluating and computing directional derivatives of the defining polynomial of a hypersurface that arises through projection. Here, we build on this approach to show that pseudo-witness sets can be used to efficiently compute gradients and Hessians (Theorem 3.8). This gives rise to an algorithm (Algorithm 4) for computing regions of the complement of hypersurfaces $\mathcal{H}$ arising through projection that circumvents the need for symbolic elimination.
Example 1.1.
An important class of hypersurfaces that are obtained through projection are discriminants of parametric polynomial systems, e.g., see [20] (see also Section 4). As a running example, we will consider one of the simplest instances of such a discriminant, the quadratic polynomial $f(a,b,z)=z^{2}+az+b$. This is a polynomial in the variables $(a,b,z)$. We view $(a,b)$ as parameters of $f$. The zero set of $f$ together with its derivative $f^{\prime}=\partial f/\partial z$ is the variety of triples $(a,b,z)$ such that $z$ is a double root of $f(a,b,z)$. These triples are zeros of the system
$F(a,b,z)=\begin{pmatrix}f(a,b,z)\\[2.15277pt] f^{\prime}(a,b,z)\end{pmatrix}=\begin{pmatrix}z^{2}+az+b\\[2.15277pt] 2z+b\end{pmatrix}$. (1.2)
The zero set of $F$ is an irreducible curve in $\mathbb{C}^{3}$. Its projection onto the parameter space $\mathbb{C}^{2}$ yields the discriminant hypersurface
${\mathcal{H}}=\bigl\{(a,b)\in{\mathbb{C}}^{2}\mid h(a,b):=a^{2}-4b=0\bigr\}.$ (1.3)
The real complement ${\mathbb{R}}^{2}\smallsetminus{\mathcal{H}}$ of the hypersurface consists of the two regions $\{(a,b)\in{\mathbb{R}}^{2}\mid h(a,b)>0\}$ and $\{(a,b)\in{\mathbb{R}}^{2}\mid h(a,b)<0\}$. Our algorithm first uses a pseudo-witness set to determine that the degree of $h$ is $\deg h=2$ and constructs the routing function
$r(a,b)=\frac{|h(a,b)|}{(1+(a-c_{1})^{2}+(b-c_{2})^{2})^{2}}.$
Using $c=(13,2)$, our algorithm finds four real critical points, and through gradient flow, it determines that three of them belong to one region, whereas the fourth belongs to another region. The result is shown in Figure 1. During that computation, our algorithm never has direct access to the polynomial $h$.
Figure 1: The quadratic discriminant $\{h(a,b)=0\}$ is the black curve in the picture. There are two regions, one above and one below the black curve. Our algorithm represents these regions by critical points (green) of a routing function (with level sets illustrated with colors between magenta and yellow). There is one critical point in the top region and three critical points in the bottom region. The latter three points are connected by gradient flow (the blue curves). The critical points and the flow trajectories were computed without direct access to the polynomial $h=a^{2}-4b$.
The remainder of this paper is organized as follows. Section 2 reviews background on routing functions, gradient roadmaps, and homotopy continuation for the case when the defining polynomial $h$ of $\mathcal{H}$ is known. Section 3 describes how to extend this approach to the case when $h$ is unknown, through the concept of pseudowitness sets, and ends with a high-level description of our method (Algorithm 4). Section 4 presents illustrative examples. The paper concludes in Section 5.
2 The case of a known defining polynomial
The following collects background information regarding how to compute the real complements of a hypersurface $\mathcal{H}=V(h)$ with known defining polynomial $h$. In particular, we recall the concept of routing functions (Section 2.1) and gradient roadmaps (Section 2.2), as well as homotopy continuation and monodromy solving (Section 2.3).
2.1 Routing points and routing functions on ${\mathbb{R}}^{n}$
This section is based on [17, Sections 2 and 3]. Let $h\in\mathbb{R}[x_{1},\ldots,x_{n}]$ be a polynomial defining a hypersurface ${\mathcal{H}}=V(h)$. The complement ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$ consists of connected components, that we call regions, on which the sign of $h$ is constant. To compute these regions, we use the routing function
$r(x)=\frac{|h(x)|}{q(x)^{e}},\quad q(x)=1+\sum_{i=1}^{n}(x_{i}-c_{i})^{2},$ (2.1)
where $c=(c_{1},\dots,c_{n})\in{\mathbb{R}}^{n}$ is a generic point and $e$ is a positive integer satisfying $2e>\deg h$. This is a rational function that behaves like $h$ near $\mathcal{H}$} but is bounded on ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$. As discussed in Section 1, the idea is to compute the critical points of $r$.
Definition 2.1.
The routing points of the routing function $r$ from (2.1) are given by the set
$Z(r)\vcentcolon=\bigl\{y\in{\mathbb{R}}^{n}\mid r(y)\neq 0,\ \nabla r(y)=0\bigr\}.$
A routing point $y\in Z(r)$ is called nondegenerate if the Hessian matrix ${\rm hess}(r)(y)$ is invertible. The index of a nondegenerate routing point $y$ is the number of positive eigenvalues of ${\rm hess}(r)(y)$. Non-zero eigenvectors of ${\rm hess}(r)(y)$ to positive eigenvalues are called unstable eigenvector directions.
The next theorem summarizes the properties of the routing function $r$ from [17] that allow us to use it for computing the connected components of ${\mathbb{R}}^{n}\smallsetminus\mathcal{H}$.
Theorem 2.2 ([17, Theorem 2.8, Proposition 3.2]).
Let $r$ be as in (2.1), with $e>\deg(h)/2$. Then, there exists a nonempty Zariski open subset $\mathcal{U}\subset\mathbb{R}^{n}$ such that for every $c=(c_{1},\ldots,c_{n})\in\mathcal{U}$, the following properties hold:
  1. 1.
    For all $\epsilon>0$, there exists $\delta>0$ such that if $x\in{\mathbb{R}}^{n}$ with $\lVert x\rVert\geq\delta$, then $\lvert r(x)\rvert<\epsilon$.
  2. 2.
    The set $Z(r)$ of routing points of $r$ is finite and each routing point is nondegenerate.
  3. 3.
    For each $\alpha\in{\mathbb{R}}_{>0}$, there is at most one $y\in Z(r)$ with $r(y)=\alpha$.
  4. 4.
    The norms of $r$, $\nabla r$, and ${\rm hess}(r)$ are bounded on ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$.
  5. 5.
    Each region of ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$ contains at least one routing point of $r$ of index 0.
Example 2.3.
We continue our running example. Recall from Example 1.1 the discriminant polynomial $h(a,b)=a^{2}-4b$. Taking $e=2$ and $(c_{1},c_{2})=(13,2)$, we obtain
$r(a,b)=\frac{|a^{2}-4b|}{\big(1+(a-13)^{2}+(b-2)^{2}\big)^{2}},$
which vanishes at infinity and satisfies the conditions of Theorem 2.2. Thus $r$ is a routing function on ${\mathbb{R}}^{2}$ and it has four (nondegenerate) routing points approximately at
$(-12.339,-2.107),\quad(-3.918,-6.636),\quad(13.040,1.994),\quad\text{and}\quad(3.217,8.083)$
with indices 0, 1, 0, and 0, respectively. Each region of ${\mathbb{R}}^{2}\smallsetminus{\mathcal{H}}$ contains a routing point of index 0 as guaranteed by Theorem 2.2; see Figure 1.
2.2 Gradient flow and connectivity via routing functions
A routing function $r$ together with its routing points $Z(r)$ from Section 2.1 constitute a powerful numerical representation of the complement ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$. The properties of $r$ allow us to determine which of the routing points belong to the same region. This, in turn, yields an efficient membership test for each of the regions as well as the Euler characteristic of each region.
The key to connecting the routing points that belong to the same region is the Mountain Pass Theorem [1, 2, 43], which (in our setting) implies that all routing points of index 0 can be connected via gradient flow emanating from routing points of higher index [17, Proposition 4.2]. This gives rise to Algorithm 1, which is an adaptation of [17, Algorithm 1]. The algorithm is illustrated by Figure 1, where the three routing points that belong to the same region are connected via a single routing point of index 1.
Input: A routing function $r$ on ${\mathbb{R}}^{n}$ associated with a defining polynomial $h$
Output: A partition $\{C_{1},\dots,C_{s}\}$ of the set of routing points $Z(r)$ of $r$ on ${\mathbb{R}}^{n}$ corresponding to the connected components of ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$, where ${\mathcal{H}}=V(h)$.
1ex(1) Find routing points and their indices. Compute all routing points $Z(r)=\{y_{1},\dots,y_{m}\}\subset{\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$ of $r$ and their corresponding indices $i_{1},\dots,i_{m}$.
(2) Initialize adjacency matrix. Set $A=I_{m}$, the $m\times m$ identity matrix.
(3) Populate adjacency matrix via unstable directions.
for $j=1,\dots,m$ do
 foreach unstable eigenvector direction $v$ for ${\rm hess}(r)(y_{j})$ do
      Compute limit routing point from $y_{j}$ in the direction $v$ with respect to $r$, say $y_{w_{+}}$.
      Set $A_{jw_{+}}=A_{w_{+}j}=1$.
      Compute limit routing point from $y_{j}$ in the direction v-v with respect to $r$, say $y_{w_{-}}$.
      Set $A_{jw_{-}}=A_{w_{-}j}=1$.
   end foreach
end for
1ex(4) Transitive closure. Set $M$ to be the transitive closure of $A$.
Partition $Z(r)$ based on the connected components of $M$, say $C_{1},\dots,C_{s}$.
(5) Return $\{C_{1},\dots,C_{s}\}$.
Algorithm 1 Connected components
Once we have partitioned the routing points into subsets $C_{1},\ldots,C_{s}\subset Z(r)$ as in Algorithm 1, we can use each $C_{i}$ as a representation of the region of ${\mathbb{R}}^{n}\smallsetminus\mathcal{H}$ that its routing points belong to. For instance, it can be used for membership testing: any point $z\in{\mathbb{R}}^{n}\smallsetminus\mathcal{H}$ that is not a routing point belongs to the region represented by $C_{i}$ if and only if gradient flow with respect to $r$ from $z$ gives a path that converges to a routing point in $C_{i}$ [17, Proposition 4.1]. Furthermore, the Euler characteristic of the region represented by $C_{i}$ is given by $\sum_{j=0}^{n}(-1)^{j}\mu_{j}$, where $\mu_{j}$ is the number of routing points of index $j$ in $C_{i}$ [17, Theorem 3.8].
2.3 Homotopy continuation and monodromy solving
In the previous two sections we explained how to use a routing function $r$ for computing regions. The key is computing routing points. These are solutions to a system of rational functions $\nabla r=0$, and can be computed using homotopy continuation.
Homotopy continuation is a foundational numerical method in algebraic geometry. For a detailed introduction, we refer the reader to [5, 50]. All homotopies considered here have the form
$\begin{split}H\colon{\mathbb{C}}^{n}\times[0,1]&\longrightarrow{\mathbb{C}}^{n}\\ (x,t)&\longmapsto H(x,t),\end{split}$
where $H(x,t)$ is analytic in both $x$ and $t$. Given a start point $x^{*}\in{\mathbb{C}}^{n}$ satisfying
$H(x^{*},1)=0,\quad\det J_{x}H(x^{*},1)\neq 0,$ (2.2)
where $J_{x}H$ denotes the Jacobian of $H$ with respect to $x$, the Implicit Function Theorem guarantees a smooth local path $x\colon{\mathbb{C}}\to{\mathbb{C}}^{n}$ near $t=1$ such that $H(x(t),t)\equiv 0$ and $x(1)=x^{*}$. Following [29], this path is trackable if all points along it are nonsingular solutions of $H(\bullet,t)=0$ for $t\in(0,1]$. A trackable path can be numerically followed using a predictor–corrector scheme to integrate the Davidenko differential equation
$\dot{x}(t)=-J_{x}H(x,t)^{-1}\,J_{t}H(x,t),\quad H(x(t),t)=0,$ (2.3)
for $t\in(0,1]$. If the limit $\lim_{t\to 0}x(t)$ exists in ${\mathbb{C}}^{n}$, the path is said to converge to an endpoint $x(0)$ satisfying $H(x(0),0)=0$; otherwise, the path diverges.
Example 2.4.
Let $F(a,b,z)$ be as in Example 1.1. We wish to intersect the curve $V(F)\subseteq{\mathbb{C}}^{3}$ with the plane V(a1)V(a-1) by solving $F(1,x_{1},x_{2})=0$. For this, we can consider the homotopy
$H(x,t)=(1-t)F(1,x_{1},x_{2})+\gamma\,t\begin{pmatrix}x_{1}^{2}-1\\[2.0pt] x_{2}-1\end{pmatrix}=0,$
where $\gamma=1+\sqrt{-1}$ ensures genericity. Explicitly,
$H(x,t)=(1-t)\begin{pmatrix}x_{2}^{2}+x_{1}x_{2}+1\\[2.0pt] 2x_{2}+x_{1}\end{pmatrix}+\gamma\,t\begin{pmatrix}x_{1}^{2}-1\\[2.0pt] x_{2}-1\end{pmatrix}$.
The start points $(1,1)$ and (1,1)(-1,1) are nonsingular solutions of $H(x,1)=0$. Both corresponding paths are trackable and converge respectively to (2,1)(2,-1) and (2,1)(-2,1) of $H(x,0)=0$. This shows that the intersection of the zero set of $F$ with the plane $a=1$ is $\{(2,1,-1),(-2,1,1)\}$.
When only a subset of the solutions to a system of polynomials or rational functions is known, additional solutions can often be found through monodromy [25, 47, 19, 7, 28]. In this approach, we view our target system as a member of a parametric family of systems, and let the parameters vary continuously along a loop in parameter space, while numerically tracking a known solution. When the loop returns to its starting point, the endpoint of the tracked path is again a solution of the original system, but not necessarily the same one. This induces an action of the fundamental group of the parameter space on the fibre of solutions over a base parameter, called the monodromy action. The image of this action inside the permutation group of the solutions is the monodromy group. See [27] for further discussion of monodromy groups and their numerical computation. A transitive monodromy group implies that all (complex) solutions of the system can be reached from a single start solution via finitely many monodromy loops.
Example 2.5.
Recall that in Example 2.4, we computed the intersection of the curve $V(F)$ with the plane V(a1)V(a-1) by solving $F(1,x_{1},x_{2})=0$. We can vary that plane in a monodromy loop: For a loop $\Gamma\colon[0,1]\to{\mathbb{C}}$ with $\Gamma(0)=\Gamma(1)$, we consider now the new homotopy $H(x,t)=F(x_{1},\Gamma(t),x_{2})=0$ with start point (2,1)(2,-1). If $\Gamma(t)=3/4+(1/4)e^{2\pi i(1-t)}$ the corresponding path is trackable and returns to (2,1)(2,-1). In contrast, for the loop $\Gamma(t)=1/4+(3/4)e^{2\pi i(1-t)}$ the path ends at (2,1)(-2,1). Hence, the monodromy action exchanges these two solutions. The reason why the second loop gives a new solution, while first does not, is that the second loop encircles a so-called branch point.
In our case, we embed our target system $\nabla r(x)=0$ into the family $\nabla r(x)-q=0$ for parameters $q\in{\mathbb{C}}^{n}$. As a start solution, we take a generic point $x_{0}\in{\mathbb{C}}^{n}$, which by design is a solution of $\nabla r(x)-q_{0}=0$ if we set $q_{0}=\nabla r(x_{0})$. Since the incidence variety $\{(x,q)\mid\nabla r(x)-q=0\}$ is a graph and hence irreducible, it follows that the monodromy action is transitive [19]. Consequently, we can compute all solutions of $\nabla r(x)-q_{0}=0$ by tracking monodromy loops. The routing points are then obtained by tracking those solutions along $H(x,t)=\nabla r(x)-t\,q_{0}$ from $t=1$ to $t=0$.
2.4 Summary of the algorithm for known defining polynomial
In the subsections above, we have discussed all the ingredients needed to compute the regions of ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$ for a known polynomial $h$, and we summarize this as Algorithm 2, which is an adaptation of [17, Algorithm 1]. This is also the foundational algorithm behind the Julia package HypersurfaceRegions.jl [11]. The goal of the rest of the paper is to adapt this to the case when the defining polynomial $h$ is unknown, and the hypersurface $\mathcal{H}$ instead arises as the projection of a known variety.
Input: A nonzero polynomial $h\in\mathbb{R}[x_{1},\ldots,x_{n}]$ defining a hypersurface $\mathcal{H}$.
Output: The regions of the real complement ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$.
1ex(1) Construct a routing function. Randomly select $c\in\mathbb{R}^{n}$ and take $e>\deg(h)/2$. This defines a routing function $r(x)=|h(x)|/q(x)^{e}$ as in (2.1).
(2) Compute routing points. Compute all solutions to $\nabla r(x)=0$ through monodromy, as described in Section 2.3. Filter out the real solutions. These are the routing points.
(3) Compute regions. Run steps (2)–(5) in Algorithm 1.
Algorithm 2 Computing regions of a hypersurface with known defining polynomial
3 The case of an unknown defining polynomial
In this section, we treat the main problem of interest in the paper, namely, to compute the complement ${\mathbb{R}}^{k}\smallsetminus\mathcal{H}$ of a hypersurface $\mathcal{H}$ with unknown defining polynomial $h\in\mathbb{R}[p_{1},\ldots,p_{k}]$, which is defined as the projection of some known variety $X$. More precisely, we consider a variety $X=V(F)\subset{\mathbb{R}}^{n}$ given by a known system $F\in({\mathbb{R}}[p_{1},\ldots,p_{k},z_{1},\ldots,z_{n-k}])^{n-k+1}$ such that
$\mathcal{H}=\overline{\pi(X)},$
where $\pi\colon{\mathbb{C}}^{k}\times{\mathbb{C}}^{n-k}\to{\mathbb{C}}^{k}$ is the projection $\pi(p,z)=p$, and the line denotes Zariski closure.
In what follows, we will let $h$ be a polynomial of minimal total degree such that $\mathcal{H}=V(h)$ (this defines $h$ uniquely up to scaling).
The key observation is that in order to carry out step (1)–(3) of Algorithm 2, we do not need to have an explicit expression for $h$. Instead, we just need the following:
  • We need to know $\deg h$ in order to choose the denominator of the routing function in step (1).
  • We need to be able to evaluate $\nabla h$ and ${\rm hess}(h)$ in order to do homotopy continuation in step (2) and gradient flow in step (3).
It turns out that both these pieces of information about $h$ are available through a pseudo-witness set of $\mathcal{H}$. We explain this concept and how it gives the degree in Section 3.1. We then proceed to derive algorithms for evaluating $\nabla r$ and ${\rm hess}(r)$ through the data of a pseudo-witness in Section 3.2, which is the main theoretical contribution of the paper. Building on this, we give a strategy (Algorithm 4) for computing ${\mathbb{R}}^{k}\smallsetminus\mathcal{H}$ in Section 3.3.
Throughout this section, we will work with the logarithmized routing function
$\log(r)=\log(|h|)-e\log(q),$
which shares critical points with $r$, but is substantially more well-behaved numerically [11]. In particular, it turns the formulas in Section 3.2 into sums instead of products of high degrees.
Remark 3.1.
It is often useful to consider the complement of the union of $\mathcal{H}$} with another hypersurface $V(g)$ where $g$ is a known polynomial, e.g., when imposing positivity conditions [17, 16]. This situation can be easily handled by simply adding $\log(g)$ to $\log(r)$ above and adjusting $e$ to ensure that $2e>\deg(h)+\deg(g)$.
3.1 Witness and pseudo-witness sets
Witness and pseudo-witness sets are data structures for representing (positive-dimensional) algebraic varieties numerically. We give a brief introduction to this topic. General references for witness and pseudo-witness sets are the books [5, 50].
Let us first define the notion of a witness set. Let $F=(f_{1},\ldots,f_{r})\in({\mathbb{C}}[x_{1},\ldots,x_{n}])^{r}$ be a system of $r$ polynomials and consider the corresponding variety $V(F)\subset{\mathbb{C}}^{n}$. Suppose that $X\subset V(F)$ is a $d$-dimensional irreducible component of $V(F)$. A witness set represents $X$ by the intersection of $X$ with a general linear space $\mathcal{M}$ of codimension $d$. If $\mathcal{M}$ is sufficiently generic, it follows from Bézout’s theorem that it intersects $X$ in $\deg X$ points. Below is the full definition.
Definition 3.2.
  1. 1.
    Let $X\subset\mathbb{C}^{n}$ be an irreducible variety of dimension $d$. A witness set for $X$ is a triple $(F,\mathcal{M},\mathcal{W})$, where $F$ is a system such that $X$ is an irreducible component of $V(F)$, $\mathcal{M}\subset{\mathbb{C}}^{n}$ is a general linear subspace of codimension $d$ and $\mathcal{W}=X\cap\mathcal{M}$ is a finite set of points. Here, $F$ is called a witness system, $\mathcal{M}$ is the witness slice, and $\mathcal{W}$ is the corresponding witness point set.
  2. 2.
    We call the witness set $(F,\mathcal{M},\mathcal{W})$ reduced, if for all points $x\in\mathcal{W}$ we have $\mathrm{rank}(JF(x))=n-d$. Here, $JF(x)$ denotes the Jacobian of $F$ at $x$.
  3. 3.
    If $X$ is of pure dimension $d$, but not necessarily irreducible, let $X_{1},\ldots,X_{\ell}$ be its irreducible components. We call $(F,\mathcal{M},\mathcal{W}_{1}\cup\cdots\cup\mathcal{W}_{\ell})$ a witness set for $X$, where $\mathcal{W}_{i}=X_{i}\cap\mathcal{M}$.
It is important to note that the defining equations for $X$ are not used in Definition 3.2; instead, $X$ is represented using $F$. This can potentially introduce multiplicities, which is why we have the second item in the definition. The second item can be rephrased as follows: A witness set for an irreducible component $X$ of $V(F)$ is reduced when the ideal in the primary decomposition of $\langle F\rangle\vcentcolon=\langle f_{1},\ldots,f_{r}\rangle$ corresponding to $X$ is radical. In a reduced witness set, the points in $X\cap\mathcal{M}$ are nonsingular solutions to a square polynomial system in such a way that homotopy continuation can be used to trace the witness set to another witness set $X\cap\mathcal{M}^{\prime}$ for a different slice $\mathcal{M}^{\prime}$. This way, reduced witness sets enable several numerical operations in homotopy continuation, including sampling points on a variety, testing membership, and computing intersections [4, 31, 33, 48, 49, 50]. If a witness set $(F,\mathcal{M},{\mathcal{W}})$ is not reduced, one can utilize deflation to replace the witness system $F$ with a new witness system $F^{\prime}$ so that $(F^{\prime},\mathcal{M},{\mathcal{W}})$ is reduced [32].
Example 3.3.
We return to our running example from Example 1.1. Let $F(a,b,z)$ as in (1.2) and recall that the zero set $V(F)\subset{\mathbb{C}}^{3}$ of $F$ is an irreducible curve. Let $X=V(F)$ and intersect it with a general two-dimensional plane $\mathcal{M}$, say a+5bz10=0a+5b-z-10=0. The intersection $\mathcal{M}\cap X$ is illustrated in Figure 2. It consists of two points (in particular, this shows that $\deg X=2$). The witness set of $X$ is $(F,\mathcal{M},\mathcal{W})$ and it consists of $F$, the plane $\mathcal{M}$ and the two intersection points in $\mathcal{W}=X\cap\mathcal{M}$. The witness set $(F,\mathcal{M},\mathcal{W})$ is reduced.
[…]
Figure 2: A witness set for the red curve $X=V(F)$ with $F$ as in Example 3.3. The witness set is given by the intersection of the red curve $X$ with the green linear space $\mathcal{M}$. The intersection consists of the two points labeled $\mathcal{W}$.
Example 3.4.
We give an example of a non-reduced witness-set. Consider the case $n=2$ and the polynomial $F(x_{1},x_{2})=(x_{1}^{2}+x_{1}^{2}-1)^{2}$. The zero set $X=V(F)$ is the circle. It has dimension $d=1$. The intersection of $V(F)$ with a general line consists of $2$ points.
Take, for instance, the line $\mathcal{M}=V(x_{1})$. Then, $\mathcal{W}=X\cap\mathcal{M}=\{(0,1),(0,-1)\}$. However, $JF(x_{1},x_{2})=4(x_{1}^{2}+x_{1}^{2}-1)\,(x_{1},x_{2})$, so that $JF(0,1)=(0,0)$ and JF(0,1)=(0,0)JF(0,-1)=(0,0) both have rank 0, and not nd=21=1n-d=2-1=1. The witness set $(F,\mathcal{M},\mathcal{W})$ is therefore not reduced. This is caused by the fact that the polynomial $x_{1}^{2}+x_{1}^{2}-1$ defining the circle appears with multiplicity 2 in $F$.
Next, we consider pseudo-witness sets. These represent the Zariski closure of images of projections of $X$. For this, let $F$ be in the variables $x=(x_{1},\dots,x_{n})$, with $p:=(x_{1},\dots,x_{k})$ denoting the first $k$ variables and $z:=(x_{k+1},\ldots,x_{n})$ the last nkn-k variables. We then consider the projection
$\pi\colon{\mathbb{C}}^{n}\to{\mathbb{C}}^{k},\quad\pi(p,z)=p,$ (3.1)
and let
${\mathcal{H}}=\overline{\pi(X)}\subset{\mathbb{C}}^{k}$ (3.2)
denote the Zariski closure of the image of $X$. When $\mathcal{H}$} is a hypersurface whose defining polynomial is not explicitly known, we can still represent it numerically by a pseudo-witness set. The idea is to lift lines in $\mathbb{C}^{k}$ to a linear space in $\mathbb{C}^{n}$. This works as follows. Consider a general line $\mathcal{L}\subset{\mathbb{C}}^{k}$ and a general linear subspace $\mathcal{L}^{\prime}\subset{\mathbb{C}}^{n-k}$ of dimension nd1n-d-1. Then, the intersection of $\mathcal{L}\times\mathcal{L}^{\prime}$ with $X$ is a finite set $\mathcal{W}$ of points of size $|\mathcal{W}|=\deg{\mathcal{H}}\cdot\deg\pi^{-1}(p)$ where $p\in{\mathcal{H}}$ is general.
Projecting the coordinates of points in $\mathcal{W}\subset{\mathbb{C}}^{n}$ to the first $k$ coordinates yields $\pi(\mathcal{W})={\mathcal{H}}\cap\mathcal{L}$. In other words, we can compute the intersection of ${\mathcal{H}}\subset{\mathbb{C}}^{k}$ with a general line $\mathcal{L}$ by computing intersection of $X$ with $\mathcal{L}\times\mathcal{L}^{\prime}$. This yields the notion of pseudo-witness set, which we summarize in the following definition.
Definition 3.5.
  1. 1.
    Let $X\subset\mathbb{C}^{n}$ be an irreducible variety of dimension $d$, such that ${\mathcal{H}}=\overline{\pi(X)}$ be a hypersurface. A pseudo-witness set for $\mathcal{H}$} is a quadruple $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$, where $F$ is a system such that $X$ is an irreducible component of $V(F)$, $\mathcal{L}$ is a general line and $\mathcal{L}^{\prime}$ a general linear space $\mathcal{L}^{\prime}$ of dimension nd1n-d-1, and $\mathcal{W}=X\cap(\mathcal{L}\times\mathcal{L}^{\prime})$.
  2. 2.
    We call the pseudo-witness set $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$ reduced, if $\mathrm{rank}(JF(x))=n-d$ for all points $x\in\mathcal{W}$. Here, $JF(x)$ denotes the Jacobian of $F$ at $x$.
  3. 3.
    If $X$ is of pure dimension $d$, but not necessarily irreducible, let $X_{1},\ldots,X_{\ell}$ be its irreducible components such that $\overline{\pi(X_{j})}$ is a hypersurface. We call $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W}_{1}\cup\cdots\cup\mathcal{W}_{\ell})$ a pseudo-witness set for $\mathcal{H}=\overline{\pi(X)}$, where $\mathcal{W}_{i}=X_{i}\cap(\mathcal{L}\times\mathcal{L}^{\prime})$.
The method we describe in Section 3.2 for evaluating the gradient and the Hessian of the routing function works for any pseudo-witness set. However, as discussed above, once we want to use homotopy continuation and track a pseudo-witness set from one linear space $\mathcal{L}\times\mathcal{L}^{\prime}$ to another $\hat{\mathcal{L}}\times\hat{\mathcal{L}}^{\prime}$ we need to work with reduced pseudo-witness sets, so that the second condition in (2.2) is satisfied. See also Remark 3.11 below.
We note that Definition 3.5 extends to varieties of arbitrary codimension, not just hypersurfaces. Here, we only need the hypersurface case.
Example 3.6.
We continue Examples 1.1 and 3.3. Recall that the discriminant $\mathcal{H}$} is defined by the polynomial $h(a,b)=a^{2}-4b$. It is obtained after eliminating $z$ from $F(a,b,z)$, where $F$ is as in (1.2). Consider the line $\mathcal{L}=\{p+tv\mid t\in\mathbb{C}\},$ where $p=(0,2)$ and v=(2,3/5)v=(-2,3/5). We compute the intersection $\mathcal{H}$$\cap\mathcal{L}$ using pseudo-witness sets for $X=V(F)$. Since $X$ is a curve ($d=1$) and since $n=3$, we must take a general linear space $\mathcal{L}^{\prime}$ of dimension nd1=311=1n-d-1=3-1-1=1 in $\mathbb{C}$. There is only one option, namely $\mathcal{L}^{\prime}=\mathbb{C}$. Computing the intersection of $X$ with $\mathcal{L}\times\mathcal{L}^{\prime}$ yields
$\mathcal{W}=X\cap(\mathcal{L}\times{\mathbb{C}})=\Bigl\{\Bigl(-\tfrac{3+\sqrt{209}}{5},\tfrac{109+3\sqrt{209}}{50},\tfrac{3+\sqrt{209}}{10}\Bigr),\Bigl(-\tfrac{3-\sqrt{209}}{5},\tfrac{109-3\sqrt{209}}{50},\tfrac{3-\sqrt{209}}{10}\Bigr)\Bigr\}.$
Since $\pi(\mathcal{W})={\mathcal{H}}\cap\mathcal{L}$, the intersection of the discriminant with $\mathcal{L}$ is given by taking the first two coordinates of these points. The pseudo-witness set $(X,\pi,\mathcal{L}\times{\mathbb{C}},\mathcal{W})$ is reduced.
[…]
Figure 3: The left picture illustrates a pseudo-witness set for the red curve $X=V(F)$. While Figure 2 the green linear space $\mathcal{M}$ was general, here it must be a product space: $\mathcal{M}=\mathcal{L}\times\mathbb{C}$, where $\mathcal{L}$ is a general line; see Example 3.6. The pseudo-witness set is then given by the intersection of the red curve $X$ with $\mathcal{L}\times\mathbb{C}$. The projection of $\mathcal{W}$ onto the yellow two-dimensional plane yields $\mathcal{H}$$\cap\mathcal{L}$. The right picture shows the situation within the yellow plane.
3.2 Evaluating the gradient and the Hessian
We now turn to the problem of evaluating $\nabla r$ and ${\rm hess}(r)$ of the unknown defining polynomial $h$ of $\mathcal{H}\subset{\mathbb{C}}^{k}$, using the data of a pseudo-witness set $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$ where $\mathcal{L}\subset{\mathbb{C}}^{k}$ is a line. We underline that the pseudo-witness set need not be reduced here. The pseudo-witness set in this section is only used to give access to the points in $\mathcal{H}\cap\mathcal{L}$.
It is described in [22] how to utilize a pseudo-witness set to evaluate and differentiate $h$ along the line $\mathcal{L}$. This implies that $\nabla h(p)$ can be computed from $k$ linearly independent lines $\mathcal{L}$ through $p$. Likewise, ${\rm hess}(h)(p)$ can be computed through $k^{2}$ lines through $p$. In this section, we prove that, in fact, a single general line $\mathcal{L}$ through $p$ suffices to determine both $\nabla h(p)$ and ${\rm hess}(h)(p)$. This drastically improves the complexity of the computation.
Let $\mathcal{L}\subset{\mathbb{C}}^{k}$ be a general line that passes through a given point $p\in{\mathbb{C}}^{k}$ in the direction $b$, and consider a general linear space $\mathcal{L}^{\prime}$ of dimension nd1n-d-1. Recall from Section 3.1 that under the assumption of sufficient genericity, the line $\mathcal{L}$ meets $\mathcal{H}$} transversely in $\deg h$ points and no intersection occurs at infinity. Let
${\mathcal{H}}\cap\mathcal{L}=\{p+t_{j}b\mid j=1,\ldots,\deg h\}.$
Then, we have that
$h(p+tb)=C(b)\cdot(t-t_{1})\cdots(t-t_{\deg h})$ (3.3)
for some constant $C(b)$ that depends on $b$. In fact, for generic $b$, $C(b)\neq 0$ and $C(b)\cdot t^{\deg h}$ is the leading (highest-total degree) part of $h(p+tb)$. After logarithmizing, we obtain
$\log|h(p)|=\log|C(b)|+\sum_{j=1}^{\deg h}\log|t_{j}|.$ (3.4)
Thus, by keeping $b$ constant, we can evaluate $\log|h(p)|$ up to an unknown constant $\log|C(b)|$.
A key component in what follows next is the gradient of the $t_{i}$ with respect to $p$ and $b$. For computing these gradients, we choose parameterizations
$\mathcal{L}(t)=p+tb\quad\text{and}\quad\mathcal{L}^{\prime}(y)=v+Ay,$
where $p\in{\mathbb{C}}^{k}\smallsetminus{\mathcal{H}}$ and $b\in{\mathbb{C}}^{k}\smallsetminus\{0\}$, and $A\in\mathbb{C}^{n\times(n-d-1)}$ has full rank and $v\in{\mathbb{C}}^{n}$.
Since $p\notin{\mathcal{H}}$, it will be helpful to work with the inverse $s\coloneq 1/t$ instead of $t$. A change of variables from $t$ to $s$ yields the polynomial system
$G(s,y)=F(p+(1/s)b,v+Ay).$
For simplicity, we also denote the projection $\pi(s,y)=s$. The zeros of $G$ are given by
$V(G)=\bigcup_{j=1}^{\deg h}\pi^{-1}(s_{j}),\quad\text{ where }s_{j}:=\frac{1}{t_{j}}.$
For every $j$, pick a fixed $y_{j}\in\pi^{-1}(s_{j})$. We keep $A$ and $v$ fixed and interpret $p,b$ as variables. The Implicit Function Theorem implies that locally around every $(s_{j},y_{j})$ there exists smooth functions
$s_{j}(p,b)\quad\text{and}\quad y_{j}(p,b),\qquad j=1,\ldots,\deg h,$ (3.5)
such that $G(s_{j}(p,b),y_{j}(p,b))=0$. Differentiating that equation with respect to $p$ and $b$ yields
$J_{(s,y)}G\begin{pmatrix}\nabla_{p}s\\ J_{p}y\end{pmatrix}+J_{p}G$ $=0,$ (3.6)
$J_{(s,y)}G\begin{pmatrix}\nabla_{b}s\\ J_{b}y\end{pmatrix}+J_{b}G$ $=0.$
Differentiating the first system in (3.6) again with respect to each coordinate $b_{i}$ yields
$J_{(s,y)}G\begin{pmatrix}\partial_{b_{i}}(\nabla_{p}s)\\ \partial_{b_{i}}(J_{p}y)\end{pmatrix}+R^{(i)}(p,b)=0,$ (3.7)
where $R^{(i)}(p,b)$ depends only on $\nabla_{p}s,J_{p}y,\nabla_{b}s,J_{b}y$ and derivatives of the Jacobian matrices. Solving (3.7) yields $\nabla_{b}(\nabla_{p}s)$.
The following gives an expression of the gradient of $\log|h|$ along $\mathcal{L}$. This expression will be used to obtain formulas for the gradient and Hessian of $\log|h|$.
Proposition 3.7.
The directional derivative of $\log|h|$ along $\mathcal{L}=\{p+tb\}$ is given by
$\nabla_{p}\log|h(p)|\cdot b=-\sum_{j=1}^{\deg h}s_{j}.$
Proof.
From (3.3), it follows that
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\Bigl(\log|h(p+tb)|\Bigr)\Big|_{t=0}&=\frac{\mathrm{d}}{\mathrm{d}t}\Biggl(\sum\limits_{j=1}^{\deg h}\log|t-t_{j}|\Biggr)\Biggr|_{t=0}=\sum\limits_{j=1}^{\deg h}\frac{\mathrm{d}}{\mathrm{d}t}\Big(|t-t_{j}|\Big)\Bigr|_{t=0}=-\sum_{j=1}^{\deg h}\frac{1}{t_{j}}.\end{split}$
Substituting $s_{j}=1/t_{j}$ gives the result. ∎
Taking the gradient of $\nabla_{p}\log|h(p)|\cdot b$ with respect to $b$ then gives $\nabla_{p}\log|h(p)|$. Furthermore, the gradient of $\nabla_{p}\log|h(p)|$ with respect to $p$ is the Hessian. This proves the following result.
Theorem 3.8.
The gradient and Hessian of $\log|h|$ are, respectively, given by
$\nabla_{p}\log|h(p)|=-\sum_{j=1}^{\deg h}\nabla_{b}\,s_{j}(p,b)\qquad\text{and}\qquad{\rm hess}(\log|h(p)|)=-\sum_{j=1}^{\deg h}\nabla_{b}\bigl(\nabla_{p}s_{j}(p,b)\bigr).$
The first order derivatives of $s_{j}$ are obtained by solving the linear systems in (3.6). The second order derivatives are obtained by solving the linear system (3.7). We summarize the strategy to evaluate the gradient and Hessian of the function $\log|h|$ in Algorithm 3.
Input: A pseudo-witness set $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$ for a hypersurface ${\mathcal{H}}\subset{\mathbb{C}}^{k}$ as in Section 3.1, defined by an unknown nonzero polynomial $h\in\mathbb{R}[p_{1},\ldots,p_{k}]$. A point $p\in{\mathbb{C}}^{k}$.
Output: The gradient $\nabla\log|h|$ and the Hessian ${\rm hess}(\log|h|)$ evaluated at $p$.
1ex(1) Intersect $\mathcal{H}$} with a general line $\mathcal{L}\subset{\mathbb{C}}^{k}$ through $p$. Pick random direction $b\in{\mathbb{C}}^{k}$, set $\mathcal{L}=\{p+tb\mid t\in{\mathbb{C}}\}$, and let $s_{j}(p,b)$ and $y_{j}(p,b)$ for $j=1,\ldots,\deg h$ be as in (3.5).
(2) Solve linear systems. For each $j=1,\ldots,\deg h$, compute $\nabla_{p}s_{j}$, $J_{p}y_{j}$, $\nabla_{b}s_{j}$, and $J_{b}y_{j}$ by solving the linear system (3.6), and compute $\nabla_{b}(\nabla_{p}s_{j})$ by solving (3.7).
(3) Compute the gradient and Hessian. Use the formulas from Theorem 3.8.
Algorithm 3 Evaluating the gradient and the Hessian of $\log r$
Example 3.9.
We consider again the quadratic discriminant from Example 1.1. As in Example 3.6 we consider a general line $\mathcal{L}\subset\mathbb{C}^{2}$ and $\mathcal{L}^{\prime}=\mathbb{C}$, so that $\mathcal{L}\times\mathcal{L}^{\prime}=\mathcal{L}\times$$\mathbb{C}$}. Let $p=(p_{1},p_{2})\in{\mathbb{R}}^{2}\smallsetminus{\mathcal{H}}$ and $b=(b_{1},b_{2})\in{\mathbb{R}}^{2}\smallsetminus\{(0,0)\}$ be a direction vector, so that $\mathcal{L}=\{p+tb\}$. Plugging into (1.2) we then have
$G(s,y)=F\bigl(p+(1/s)b,y\bigr)=\begin{pmatrix}y^{2}+(p_{1}+(1/s)b_{1})y+(p_{2}+(1/s)b_{2})\\ 2y+(p_{1}+(1/s)b_{1})\end{pmatrix}=0.$ (3.8)
For fixed $(p,b)$, this system has two solutions corresponding to points of $\mathcal{H}$$\cap\mathcal{L}$.
We use the strategy laid out in Theorem 3.8 to compute the Hessian of $\log|h|$. We differentiate (3.8) with respect to $p$ and $b$ to obtain equations (3.6) with Jacobian matrices
$J_{(s,y)}G=\begin{pmatrix}(-1/s^{2})(b_{1}y+b_{2})&2y+p_{1}+(1/s)b_{1}\\ (-1/s^{2})b_{1}&2\end{pmatrix},\quad J_{p}G=\begin{pmatrix}y&1\\ 1&0\\ \end{pmatrix},\quad\text{and}\quad J_{b}G=\begin{pmatrix}y/s&1/s\\ 1/s&0\\ \end{pmatrix}$.
We evaluate these matrices at the zeros of $G$ to determine $\nabla_{p}s_{j}$ and $\nabla_{b}s_{j}$.
Next, computing the systems (3.7) for $i=1,2$ and concatenating them into one system yields
$J_{(s,y)}G\begin{pmatrix}\partial_{b_{1}}(\nabla_{p}s)&\partial_{b_{2}}(\nabla_{p}s)\\ \partial_{b_{1}}(\nabla_{p}y)&\partial_{b_{2}}(\nabla_{p}y)\end{pmatrix}=-\begin{pmatrix}R^{(1)}(p,b)&R^{(2)}(p,b)\end{pmatrix}$
where
$R^{(1)}(p,b)=\frac{1}{s^{3}}\begin{pmatrix}R^{(1)}_{11}&R^{(1)}_{12}\\ R^{(1)}_{21}&R^{(1)}_{22}\end{pmatrix},\quad\text{ and }\quad R^{(2)}(p,b)=\frac{1}{s^{3}}\begin{pmatrix}R^{(2)}_{11}&R^{(2)}_{12}\\ R^{(2)}_{21}&R^{(2)}_{22}\end{pmatrix}$.
Solving this system yields $\nabla_{b}(\nabla_{p}s)\in{\mathbb{C}}^{2\times 2}$. Summing over all $s_{j}$ we obtain, according to the formula in Theorem 3.8, ${\rm hess}(\log|h|)(p)=-\nabla_{b}(\nabla_{p}s_{1})-\nabla_{b}(\nabla_{p}s_{2})$.
When $h$ appears in a routing function (2.1), the formulas from Theorem 3.8 are easy to extend, as shown by the following corollary.
Corollary 3.10.
For a routing function $r(p)=|h(p)|/q(p)^{e}$ with $q$ as in (2.1), it holds that
$\nabla_{p}\log r(p)$ $=-\sum_{j=1}^{\deg h}\nabla_{b}\,s_{j}(p,b)-\frac{2(p-c)}{q(p)},\quad\text{and}$
${\rm hess}(\log r(p))$ $=-\sum_{j=1}^{\deg h}\nabla_{b}\bigl(\nabla_{p}s_{j}(p,b)\bigr)+\frac{4}{q(p)^{2}}(p-c)(p-c)^{T}-\frac{2}{q(p)}I_{n}.$
3.3 Summary of the algorithm for unknown defining polynomial
We summarize our strategy for computing complements of hypersurfaces defined by pseudo-witness sets in Algorithm 4. An implementation of the algorithm in Julia, built on top of the numerical algebraic geometry package HomtopyContinuation.jl [10] will be provided in a forthcoming software paper. In the next section, we showcase the result of three early proof-of-concept calculations from this implementation.
Input: A system of polynomials $F(p,z)\in({\mathbb{R}}[p_{1},\ldots,p_{k},z_{1},\ldots,z_{n-k}])^{n-k+1}$, such that ${\mathcal{H}}=\overline{\pi(X)}$ is a hypersurface, where $X\subset V(F)$ has dimension k1k-1, and $\pi(p,z)=p$.
Output: The regions of the real complement ${\mathbb{R}}^{k}\smallsetminus{\mathcal{H}}$.
1ex
(1) Compute a pseudo-witness set. Find a pseudo-witness set $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$ of $\mathcal{H}$, and compute the degree $d$ of $\mathcal{H}$ as explained in Section 3.1.
(2) Construct a routing function. Sample $c\in\mathbb{R}^{k}$ and take $e>d/2$. This defines a routing function $r$. For the next steps in the algorithm, it is enough to be able to evaluate the gradient and the Hessian of $\log r$. Use the method in Section 3.2 for this.
(3) Compute routing points. Compute all solutions to $\nabla\log r(x)=0$ by a two step homotopy: First sample a random $q\in\mathbb{C}^{k}$ and compute the all complex solutions of $\nabla\log r(x)-q=0$ by varying $q$ in monodromy loops. Then, track those solutions along the homotopy $H(x,t)=\nabla\log r(x)-tq$ for $t$ from $1$ to 0. Filter out the real solutions. These are the routing points.
(4) Compute regions. Run steps (2)–(5) in Algorithm 1.
Algorithm 4 Computing regions of a hypersurface with unknown defining polynomial
Remark 3.11.
For Algorithm 4, we do not need to assume that pseudo-witness sets are reduced as long as we have a method to compute the points in $\mathcal{H}\cap\mathcal{L}$. However, when using numerical homotopy continuation for computing $\mathcal{H}\cap\mathcal{L}$ when evaluating $\nabla r$ and ${\rm hess}(r)$, we must assume reducedness. Suppose that $X_{1}\ldots,X_{\ell}$ are the irreducible components of $X$ such that, after reordering, the first $j$ are reduced, i.e., $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},X_{i}\cap(\mathcal{L}\times\mathcal{L}^{\prime}))$ for $1\leq i\leq j\leq\ell$ are reduced pseudo-witness sets. We can track these pseudo-witness sets towards other linear spaces using homotopy continuation, since only for reduced pseudo-witness sets the second condition in (2.2) is satisfied. Consequently, we can run Algorithm 4 only on the hypersurface $\hat{\mathcal{H}}=\pi(X_{1}\cup\cdots\cup X_{j})\subset\mathcal{H}$, but not on $\mathcal{H}$. We will see an example of this in Section 4.3. This is an inherent limitation of our algorithm when using numerical homotopy continuation, which is the only method fast enough to compute the numerous sequential intersections with linear spaces that are required for each evaluation of $\nabla\log|h|$ and ${\rm hess}(\log|h|)$ via Algorithm 3.
The computational bottleneck in Algorithm 4 is solving the system $\nabla\log r=0$ with monodromy, as each monodromy loop requires many evaluations of $\nabla\log|h|$ and ${\rm hess}(\log|h|)$. To speed up this step, one can extend the set of start solutions as follows. We sample some number of points in ${\mathbb{R}}^{k}$, and use (1) gradient flow to trace them to routing points (of index 0), which are solutions of $\nabla\log r=0$ and (2) Newton method on $\nabla\log r$, which, if it converges, return solutions to $\nabla\log r=0$ We then use a parameter homotopy to trace them to additional start solutions of $\nabla\log r-q_{0}=0$.
4 Case studies
In this section, we illustrate our method on three polynomial systems arising in applications: the Kuramoto model (Section 4.1), the 3RPR mechanism from kinematics (Section 4.2), and a model for the Allee effect in population dynamics (Section 4.3). Each example consists of a parameterized polynomial system
$G(p;z)\in({\mathbb{R}}[p_{1},\dots,p_{k},z_{1},\dots,z_{n-k}])^{n-k},$
with parameters $p=(p_{1},\ldots,p_{k})$ and variables $z=(z_{1},\ldots,z_{n-k})$. We study its discriminant variety, which is the hypersurface
$\Delta=\overline{\pi\bigl(\bigl\{(p,z)\in{\mathbb{C}}^{k}\times{\mathbb{C}}^{n-k}\mid G(p;z)=0,\>\det\bigl(J_{z}G(p;z)\bigr)=0\bigr\}\bigr)}\subset{\mathbb{C}}^{k}$ (4.1)
where $\pi\colon{\mathbb{C}}^{k}\times{\mathbb{C}}^{n-k}\to{\mathbb{C}}^{k}$ is the projection to the parameter space, and overline denotes the Zariski closure. The real complement ${\mathbb{R}}^{k}\smallsetminus\Delta$ has finitely many regions. In each of these regions, the system has a constant, finite number of roots. In what follows, we compute the regions using the method outlined in Algorithm 4.
4.1 Kuramoto model
The Kuramoto model is an oscillator model that captures synchronization phenomena [42, 40, 39]. It is described by the system of ordinary differential equations
$\frac{\mathrm{d}\theta_{i}}{\mathrm{d}t}=\omega_{i}-\frac{K}{N}\sum\limits_{j=1}^{N}a_{i,j}\sin(\theta_{i}-\theta_{j}),\qquad\text{for }i=1,\dots,N,$ (4.2)
where $K$ is the coupling strength, $N$ is the number of oscillators, $\Omega=(\omega_{1},\dots,\omega_{N})$ is the vector of frequencies, and $a_{i,j}\in\{0,1\}$ is the $(i,j)$-th element of the adjacency matrix of the coupling graph.
Fixing $\theta_{N}=0$ and using the identity $\sin(x-y)=\sin(x)\cos(y)-\sin(y)\cos(x)$ we substitute $s_{i}=\sin(\theta_{i})$ and $c_{i}=\cos(\theta_{i})$ to rewrite the remaining N1N-1 equations into polynomial form. Coupling these equations with the Pythagorean identity $s_{i}^{2}+c_{i}^{2}=1$, we obtain a system of 2(N1)2(N-1) polynomial equations in 2(N1)2(N-1) variables:
$\begin{split}0&=\omega_{i}+\frac{K}{N}\sum\limits_{j=1}^{N}a_{i,j}(s_{i}c_{j}-s_{j}c_{i})\\ 0&=s_{i}^{2}+c_{i}^{2}-1.\end{split}$ (4.3)
We study the Kuramoto model with three oscillators, where the graph is the triangle; i.e., $a_{i,j}=1$ whenever $i\neq j$ and 0 otherwise. In the formulation of (4.3), this gives rise to the polynomial system
$G(\omega;z)=\begin{pmatrix}(s_{1}c_{2}-c_{1}s_{2})+(s_{1}c_{3}-c_{1}s_{3})-3\omega_{1}\\[1.99997pt] (s_{2}c_{1}-c_{2}s_{1})+(s_{2}c_{3}-c_{2}s_{3})-3\omega_{2}\\[1.99997pt] s_{1}^{2}+c_{1}^{2}-1\\[1.99997pt] s_{2}^{2}+c_{2}^{2}-1\end{pmatrix}$ (4.4)
with parameters $\omega=(\omega_{1},\omega_{2})$ and variables $z=(c_{1},c_{2},s_{1},s_{2})$. The discriminant can be computed symbolically, and is given by a degree 12 polynomial in the parameters with 41 terms. Its real complement is known to have 9 connected components (see, e.g., [15, 26]). We recover this result via our method, without making use of the discriminant polynomial. Using a routing function with center $(0.47,0.43)$, we find $59$ complex critical points. Out of these, $24$ are real (these are the routing points), and using Algorithm 1, the routing points are grouped into 9 regions. Figure 4 shows the routing points and gradient flow connections, along with the known discriminant variety.
Figure 4: Routing points and connecting paths for the discriminant of the two-parameter version of the Kuramoto model with three oscillators (4.4). The left picture shows a zoomed out version of the right picture. The gradient flow is visible in blue. The arrows point towards index-zero routing points. We see that 77 out of 99 regions are inside the star-shaped figure. The last two regions are the inside of the ellipse and the unbounded region outside the ellipse. The pictures also display the level sets of the routing function $\log r$. The routing points and the flows were computed without access to the discriminant polynomial.
4.2 3RPR
The 3RPR mechanism is an important example in kinematics of a system that exhibits a nonsingular assembly-mode change: the mechanism can pass from one solution branch to another without crossing a singularity [8, 21, 34, 37, 38, 41, 26]. It consists of three prismatic legs with revolute joints, each anchored to the ground at one end and attached to a moving triangular platform at the other.
To model its motion, two coordinate frames are used: a fixed frame attached to the ground, where the anchor points lie at fixed positions, and a moving frame attached to the triangular platform, where the platform’s corner points remain fixed relative to each other. Each leg connects one anchor point to one platform corner. In Figure 5, the leg lengths are $\ell_{1},\ell_{2}$, and $\ell_{3}$, and we work with their squares $c_{i}=\ell_{i}^{2}$ for convenience. In the fixed frame, the anchor points are located at $(0,0)$, $(A_{2},0)$, and $(A_{3},B_{3})$ while in the moving frame, the platform corners are located at the points $P_{1}=(0,0)$, $P_{2}=(a_{2},0)$, and $P_{3}=(a_{3},b_{3})$.
[…]
Figure 5: An example of a 3RPR mechanism.
A fundamental problem is to determine all the ways a real motion of the mechanism can start and end in the same “home” configuration for a fixed set of leg lengths. The configuration variables are $(p,\phi)=(p_{1},p_{2},\phi_{1},\phi_{2})$, where $p=(p_{1},p_{2})$ represents the translation of the platform and $\phi=(\phi_{1},\phi_{2})$ is a unit circle representation of its rotation. This yields a polynomial system $G(a,A,b_{3},B,c;p,\phi)$ with variables $(p,\phi)$, parameters
$a=(a_{2},a_{3}),\quad A=(A_{2},A_{3}),\quad b_{3},\quad B_{3},\quad\text{and}\quad c=(c_{1},c_{2},c_{3})$
and polynomials
$g_{1}$ $=\phi_{1}^{2}+\phi_{2}^{2}-1$
$g_{2}$ $=p_{1}^{2}+p_{2}^{2}-2(a_{3}p_{1}+b_{3}p_{2})\phi_{1}+2(b_{3}p_{1}-a_{3}p_{2})\phi_{2}+a_{3}^{2}+b_{3}^{2}-c_{1}$
$g_{3}$ $=p_{1}^{2}+p_{2}^{2}-2A_{2}p_{1}+2((a_{2}-a_{3})p_{1}-b_{3}p_{2}+A_{2}a_{3}-A_{2}a_{2})\phi_{1}$
$\qquad+~2(b_{3}p_{1}+(a_{2}-a_{3})p_{2}-A_{2}b_{3})\phi_{2}+(a_{2}-a_{3})^{2}+b_{3}^{2}+A_{2}^{2}-c_{2}$
$g_{4}$ $=p_{1}^{2}+p_{2}^{2}-2(A_{3}p_{1}+B_{3}p_{2})+A_{3}^{2}+B_{3}^{2}-c_{3}.$
Following a scaled-version of the formulation in [37] and [26, Section 4], we consider a two-parameter version of the problem obtained by viewing $(c_{1},c_{2})$ as parameters and fixing
$a_{2}=1.4,\quad a_{3}=0.7,\quad b_{3}=1.0,\quad A_{2}=1.6,\quad A_{3}=0.9,\quad B_{3}=0.6,\quad c_{3}=1.0\,.$ (4.5)
In this case, it is known that the discriminant is a degree 12 curve, and that its real complement has 8 connected components [26, Section 4]. We find all of them with our method, as illustrated in Figure 6 (which can be compared with [26, Fig. 8]). More precisely, using a routing function with center $(4.72,4.33)$, we find 24 routing points, of which 15 have index zero.
Figure 6: Routing points and connecting paths for the two-parameter version of the 3RPR system. The left picture shows a zoomed out version of the right picture. The gradient flow is visible in blue. The arrows point towards index-zero routing points. The pictures also display the level sets of the routing function $\log r$. The routing points and the flows were computed without access to the discriminant polynomial.
Next, we consider a three-parameter version of the problem, where we view $(c_{1},c_{2},c_{3})$ as free parameters, and fix the remaining ones according to (4.5). In this case, symbolic elimination shows that the discriminant surface is given by a polynomial of degree 12 with 455 terms. Using a routing function centered at $(4.72,4.33,1.70)$, our method finds 60 routing points and 9 regions of the complement of the discriminant. In these regions, the real root counts 2, 4 and 6 are recorded and certified. This is in agreement with the result obtained by analyzing the symbolic polynomial with HypersurfaceRegions.jl [11].
Finally, we consider a three-parameter version of the problem where we instead view $(c_{1},c_{2},A_{2})$ as free parameters, and fix other parameters as in (4.5). In this case, we have not been able to find a symbolic expression for the discriminant surface, but its degree can readily be determined to 24 through a pseudo witness set. This example is at the limit of what our current implementation can handle, as the monodromy step did not terminate after 14 days [^1] Computations were run on a 4-socket Intel Xeon Gold 6128 system (24 cores total, 3.4 GHz) with 3 TB RAM.. At this point, 457 complex critical points for the routing function centered at $(4.72,4.33,1.70)$ had been found, of which 90 points were real. While likely not being a complete set of routing points, these points still provide a valuable sample of the parameter space. In particular, we can use them to verify that there are open regions in parameter space with the real root counts 2, 4 and 6, respectively.
4.3 Allee effect
In the study of population dynamics, the Allee effect refers to the phenomenon where a population has a higher growth rate at higher densities. There are many possible factors that can cause this effect, and it has been widely studied (see, e.g., [6]). We consider the polynomial model from [44], where the Allee effect for three patches $z=(z_{1},z_{2},z_{3})$ is modeled by the dynamical system
$\frac{\mathrm{d}z}{\mathrm{d}t}=G(a,b,z)\quad\text{ where }\quad G(a,b,z)=\begin{pmatrix}z_{1}(1-z_{1})(z_{1}-b)+a(z_{2}-z_{1})+a(z_{3}-z_{1})\\[2.5pt] z_{2}(1-z_{2})(z_{2}-b)+a(z_{1}-z_{2})+a(z_{3}-z_{2})\\[2.5pt] z_{3}(1-z_{3})(z_{3}-b)+a(z_{1}-z_{3})+a(z_{2}-z_{3})\end{pmatrix}$.
As discussed in [44], a patch can be interpreted as a habitat, a cellular compartment, or a microplate. The parameter $a>0$ models the dispersal rate between patches, and $0<b<\frac{1}{2}$ is the Allee threshold.
The steady states of this dynamical system are the solutions of the system of equations
$G(a,b,z)=0$
with parameters $p=(a,b)$ and variables $z=(z_{1},z_{2},z_{3})$. We are interested in exploring the possible number of nonnegative solutions in the biologically relevant parameter space $(0,\infty)\times(0,\frac{1}{2})$.
To this end, we extend the discriminant $\Delta$ from (4.1) to $\Delta\cup V(ab(b-\frac{1}{2}))$. As explained in Remark 3.1, on the level of routing functions, this means adding $\log(|a|)$, $\log(|b|)$ and $\log(|b-\frac{1}{2}|)$ to $\log(r)$. It is clear from inspection of $G$ that $z=0$ is always a solution for biologically relevant parameters, and that all other nonnegative solutions have strictly positive coordinates. Hence, the number of nonnegative solutions is constant in each region of $((0,\infty)\times(0,\frac{1}{2}))\smallsetminus(\Delta\cup V(ab(b-\frac{1}{2}))$.
Computing a defining equation for $\Delta$ is highly challenging [45, 51, 52], but the relevant regions have nevertheless been determined through cylindrical algebraic decomposition in [44]. We demonstrate that the regions can also be computed using our method.
An interesting property of this example is that $\Delta$ is not irreducible, and that some of its irreducible components have a non-reduced pseduo-witness set with respect to the equations in  (4.1). Hence, if these equations are used, our implementation effectively works with the subvariety $\hat{\Delta}\subset\Delta$ corresponding to the reduced part (see Remark 3.11). We can see the effect of this in the left part of Figure 7, which shows the resulting routing function, the biologically relevant routing points and the number of nonnegative solutions of $G(a,b,z)=0$ inside the green critical points. As we can see, gradient flow connects a parameter pair $(a,b)$, such that $G(a,b,z)=0$ has 15 positive solutions with another pair, for which there are 9 positive solutions.
Figure 7: Contours and critical points of a routing functions for the modified discriminant $\Delta\cup V(ab(b-\frac{1}{2}))$ of the Allee system from Section 4.3. The left figure corresponds to using the formulation (4.1), which gives the subvariety $\hat{\Delta}$. The right figure uses the radical and gives the full discriminant $\Delta$. The routing points are in green and are labeled by the number of nonnegative solutions of $G(a,b,z)=0$. In both figures, there is a routing point of the unbounded region which is outside the picture, with nonnegative root count 3. Contours (up to a constant), critical points and gradient flow were computed without access to the discriminant polynomial.
To recover the missing part of the discriminant, we could, in principle, use deflation techniques as mentioned in Remark 3.11. In this case, however, it turns out that the radical $\sqrt{I}$ of the ideal
$I=\left\langle G(a,b,z),\ \tfrac{\partial G}{\partial z}(a,b,z)\right\rangle$
can readily be computed symbolically in the computer algebra system OSCAR [18]. By using a generating set of $\sqrt{I}$, together with the standard “squaring up” technique [4, Section 2.5] for dealing with overdetermined systems, we obtain a reduced pseudo-witness set for $\Delta$. The resulting routing function and the biologically relevant routing points are shown in the right part of Figure 7, which agrees with the regions displayed in [44, Figure 3].
5 Conclusion
Since computing a defining polynomial for a hypersurface arising as the projection of another variety is a computationally challenging elimination problem, our algorithm (Algorithm 4) provides a novel method to recover the regions of the real complement of a hypersurface without computing its defining polynomial. This allows one to compute the regions of the complement even when the hypersurface arises as the projection of an algebraic set. By intersecting this hypersurface with a line, we obtain the needed numerical information to effectively perform “elimination without eliminating” as illustrated with several examples.
Acknowledgments
We would like to thank the Fields Institute for Research in Mathematical Sciences and the organizers of the Workshop on the Applications of Commutative Algebra where the initial steps of this research were conducted. We also thank Hannah Friedman for helpful discussions in the early phases of the project.
Funding
PB was supported by DFG, German Research Foundation – Projektnummer 445466444. JC acknowledges the support of the National Science Foundation Grant DMS-2402199. AKE was partially supported by National Science Foundation Grant DMS-2023239. JDH was partially supported by the National Science Foundation Grant CCF-2331400, Simons Foundation SFM-00005696, and the Robert and Sara Lumpkins Collegiate Professorship. OH was partially funded by the European Union (Grant Agreement no. 101044561, POSALG). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Research Council (ERC). Neither the European Union nor ERC can be held responsible for them. JLG was partially supported by the Robert and Sara Lumpkins Collegiate Professorship.